# ACEGEN: Reinforcement learning of generative chemical agents for drug discovery

### v2

@@ 3.2 Benchmarking RL performance for practical drug discovery
RL optimizes the policy for the cumulative future rewards as provided by either one scoring function or a sum of multiple scoring functions. However, in practical drug discovery, it is not straightforward to write the precise scoring functions that are needed and how to weigh them into a single scalar. For example, scoring functions can sometimes not produce the desired chemistry, or show clear exploitation loopholes 45, 46. In this setting, a sufficient element of regularization to a prior policy is useful, as the prior policy has learned a chemical space distribution based on a specified training dataset of choice which acts as a representation of desirable chemistry by examples. It is for this reason that methods like REINVENT have been designed heuristically to contain terms that try to enforce regularization instead of entirely using the reward signal.
@@
To account for this, we make a modification to the average and AUC of the top 10 molecules to ensure they are diverse by an ECFP4 Tanimoto similarity of less than 0.35 to each other. Moreover, we use sphere exclusion diversity (SEDiv@1k) of de novo molecules 44 which measures the proportion of molecules needed to describe chemical space in a random sample of 1,000 molecules, as a proxy for exploration. This is more representative than the number of unique molecules that could all reside in a close area of chemical space. Lastly, we remove de novo molecules that do not pass a series of filters. This includes a basic chemistry filter (B-CF): logP less than or equal to 4.5, rotatable bond count less than or equal to 7, molecular weight in the range 150 to 650 Da, only contain atoms belonging to the following set $A\in\{C,S,O,N,H,F,Cl,Br\}$, and do not violate the substructure alerts as described in 47. Moreover on the assumption that the training dataset describes desirable chemistry by examples, a target chemistry filter (T-CF): logP and molecular weight within $\mu\pm 4\sigma$ of the training dataset distribution, as well as, removing any molecule that comprises $>10\%$ novel atomic environment bits with respect to the reference molecules, as measured by ECFP4 bits (an example of this is shown in Figure 8). The combination of the basic and target chemistry filter is denoted (B&T-CF).
@@
Re-evaluating RL performance with these new metrics (Table 2) it can be seen that a smaller proportion of molecules pass the chemistry filters for REINVENT-MolOpt, A2C, PPO, and PPOD, as expected due to decreased regularization terms to optimize efficiency. Despite this, REINVENT-MolOpt achieves the highest score for the average and AUC of the top 10 diverse molecules followed by PPOD. As a proxy for exploration, REINVENT contains the highest sphere exclusion diversity followed by AHC. Interestingly, REINFORCE maintains a high proportion of molecules passing the chemistry filters despite no explicit regularization term, outperforming REINVENT in absolute performance and efficiency. Table 3 shows the alternative approach of explicitly including these requirements in the reward signal by the addition of the chemistry filters and a diversity filter to each objective. As expected, including these requirements explicitly in the reward function increases the proportion of molecules passing these filters. However, every algorithm drops in performance as measured by the average and AUC top 10 diverse molecules that pass the filters. In this case showing that better overall score optimization and efficiency (including those that measure the quality of chemistry) is achieved by implicit regularization to the prior policy. This could be a result of the specific implementation of the chemistry filters in the reward resulting in a score of 0 if they do not pass; however, optimal implementation of reward signal can in itself be considered an art. Therefore, simpler reward signals are usually better, and, achieving desirable chemistry by regularization enables the design of a simpler reward signal. The fine-tuning of the optimization algorithm, reward signals and regularization might be target-dependent and goes beyond the scope of this work, but it is enabled by ACEGEN.
@@
Valid 21.77 21.79 21.74 21.45 21.52 22.00 21.95 21.65 21.71 21.05
Top-10 Avg 15.85 15.67 17.43 16.09 17.10 14.76 14.54 16.17 14.86 15.18
Top-10 AUC 13.67 13.55 15.65 13.91 15.80 12.94 12.78 14.57 13.00 14.01
Unique 22.28 22.67 13.68 22.68 10.21 22.44 22.83 16.63 22.80 14.13
B&T-CF 14.34 14.70 7.00 13.77 5.71 18.44 18.59 13.62 17.72 11.55
B&T-CF Top-10 Avg (Div) 14.95 14.79 16.06 15.25 15.78 14.51 14.33 15.92 14.68 14.86
B&T-CF Top-10 AUC (Div) 12.91 12.81 14.61 13.11 14.60 12.79 12.66 14.35 12.89 13.74
B&T-CF Diversity (SEDiv@1k) 17.39 18.23 10.10 17.55 7.96 16.23 17.65 7.97 17.33 7.92
Table 3: Comparison of algorithm performance with or without explicit chemistry filters included in the reward signal as a scoring function. MolOpt is the normal benchmark and MolOpt-CF is the benchmark with chemistry filters and a diversity filter applied to each objective.
@@ 3.3 Ablation study of the REINVENT algorithm
One of the most popular RL algorithms used for the fine-tuning of CLMs is REINVENT 4. From a theoretical RL perspective, this algorithm can be viewed a combination of the REINFORCE algorithm, experience replay, a reward shaping and a sequence likelihood penalty term for the loss function. However, the impact of each specific mechanism is unclear. Therefore, using the modular components of ACEGEN, we conducted an ablation study to better understand the impact of reward shaping, experience replay, and sequence likelihood penalty, which are key components of the REINVENT algorithm. We used default REINVENT hyperparameters and measured performance on the MolOpt benchmark for comparison. Regarding performance, Table 4 indicates that the primary improvement comes from adding experience replay to REINFORCE. While reward shaping does improve over REINFORCE, combining it with experience replay does not further improve performance. However, adding reward shaping that links the agent policy to the prior policy does improve the chemistry quality, measured by the proportion of valid and unique molecules passing the chemistry filters, albeit with a compromise in performance. These results highlight the importance of regularization in practical applications, as discussed in the previous section. In the original REINVENT implementation, the authors also include a loss term to penalize highly likely sequences. Table 4 shows that this primarily increases exploration but negatively affects all other metrics, including reducing the proportion of molecules passing chemistry filters. Furthermore, if used without reward shaping, it results in a complete collapse in learning.
@@
Valid 21.74 21.77 21.93 21.90 21.76 21.78
Top-10 Avg 14.74 15.85 14.80 15.69 14.79 15.69
Top-10 AUC 13.21 13.67 13.21 13.57 13.11 13.60
Unique 22.35 22.28 22.46 22.29 22.80 22.63
B&T-CF 14.15 14.34 15.31 15.08 14.83 14.74
B&T-CF Top-10 Avg (Div) 13.81 14.95 13.88 14.80 13.89 14.83
B&T-CF Top-10 AUC (Div) 12.51 12.91 12.48 12.83 12.38 12.85
B&T-CF Diversity (SEDiv@1k) 17.61 17.39 17.93 17.52 18.76 18.19
Table 4: Ablation study of REINFORCE algorithm on the MolOpt benchmark with and without separate components including experience replay, reward shaping, and a high sequence likelihood penalty. Each algorithm was run 5 times with different seeds, and results were averaged and then summed over the 23 tasks. Note that all components together constitute the ACEGEN REINVENT implementation with default hyperparameters.
@@ 3.4 Case Study: De-novo generation in the 5-HT2A
To reflect more realistic situations that may arise in practical drug discovery scenarios, we apply the different RL algorithms implemented on a more challenging objective previously proposed by 39. This objective is designing 5-HT2A receptor ligands selective over the highly related D2 receptor utilizing only structural information. This is an important and relevant challenge as clinically, this off-target profile leads to undesirable extrapyramidal side-effects in the treatment of psychosis 48. To achieve this, this objective aims to minimize the docking score against 5-HT2A (PDB: 6A93), yet maximize the docking score against D2 (PDB: 6CM4), within a budget of 32,000 molecules. Both crystal structures are co-crystallized with Risperidone, reflecting the high degree of binding pocket similarity. In contrast to the proposed objective configuration, we use rDock 49 due to more permissive licensing.
@@
Table 5: Performance summary of algorithms on 5-HT2A case study. Including after applying a basic chemistry filter (B-CF) as described previously. B-CF is the fraction of molecules that pass the chemistry filter. Note that the values here are re-normalized based on the scores achieved for a subset of 5-HT2A ligands that display at least 2-fold selectivity over D2 as extracted from ChEMBL31 32. Therefore, a score of 0.91 indicates the best score observed in the known ligand subset, and a score greater than 1.0 indicates that a single molecule achieves both a better (more negative) 5-HT2A docking score and (more positive) D2 docking score than seen anywhere in the known subset.
@@ 3.5 Case study: Scaffold constrained generation
ACEGEN can conduct constrained generation by leveraging PromptSMILES 43 which proposes iterative CLM prompting on SMILES in combination with RL to achieve fine-tune a pre-trained CLMs to the task of constrained generation. Here we demonstrate this application with two experiments: 1) scaffold decoration via known synthetic reactions as proposed with LibINVENT 18, and 2) scaffold decoration followed by scaffold-constrained docking to explore growth vectors inside a binding pocket.
In the first experiment, we compared the performance of RL algorithms on two tasks proposed with LibINVENT 18, both of which conduct scaffold decoration of a piperazine core with an amine linker moiety commonly present in D2 receptor ligands. As in the MolOpt benchmark, we apply a budget of 10,000 molecules. The first task is to optimize the predicted probability of D2 activity as evaluated by a QSAR model. The second task adds a selective reaction filter, rewarding piperazine growth via a Buchwald reaction and amine linker growth via an amide coupling reaction. Table 6 shows successful optimization of both objectives, with AHC performing best on the first task, and REINVENT (either with default or MolOpt parameters) performing best on the second task. However, all algorithms except A2C are able to solve the task with a yield above 90%, and fully satisfy the selective reaction objectives in 80 to 90% of molecules.
@@
D2 Yield 0.977 0.989 0.987 0.991 0.354 0.942 0.952
Average score 0.679 0.720 0.743 0.794 0.723 0.749 0.760
D2 with Yield 0.972 0.992 0.988 0.990 0.555 0.950 0.947
reaction Average score 0.579 0.668 0.796 0.723 0.537 0.705 0.726
filters Ratio of satisfied reaction filters 0.896 0.753 0.921 0.809 0.352 0.830 0.858
Table 6: Algorithm comparison in combination with PromptSMILES for constrained molecule generation on LibINVENT DRD2 tasks (without and with selective reaction filters). Each algorithm ran 5 times with different seeds, and results were averaged.
In the second experiment, we used PromptSMILES in combination with AHC (based on the superior performance in the 5-HT2A case study) to conduct scaffold decoration in the catalytic site of BACE1, a key therapeutic target for the treatment of Alzheimer’s disease. Here we used the reference structure 4B05 co-crystallized with AZD3839 50 which is a typical BACE1 inhibitor consisting of an amidine core interacting with two catalytic aspartates, and occupation of the P2’, P1 and P3 sub-pockets 51. The bicyclic, amidine-containing core was used as a prompt for PromptSMILES, allowing two growth vectors, one into the P2’ sub-pocket, and one into the P1 and P3 sub-pockets. Molecules were rewarded by minimizing the docking score using substructure constrained docking, as well as maintaining the number of heavy atoms and rotatable bonds in a sensible range (see Appendix F). Figure 2 shows the successful maximization of the reward over the generation of 10,000 molecules (Figure 2a,b), and corresponding minimization of the docking score (Figure 2c). The top 10 molecules identified contain improved docking scores in the range -23.9 to -15.8 compared to the reference molecule re-docked with a docking score of -8.44. Docked poses (shown in Figure 15) show the recovery of aromatic rings in the P1 and P2 sub-pockets. This case study highlights the successful constrained optimization of a complex objective.
Figure 2: Optimization of the multi-objective reward. The average reward and optimization of the underlying docking score are shown. The top 10 de novo molecules are shown by multi-objective reward, with the constrained substructure highlighted in red and the docking score labeled below. For reference, the co-crystal ligand is re-docked with a docking score of -8.44. The docked poses are shown in Figure 15.

### v3

@@ 3.2 Benchmarking RL performance for practical drug discovery
RL optimizes the policy for the cumulative future rewards as provided by either one scoring function or a sum of multiple scoring functions. However, in practical drug discovery, it is not straightforward to write the precise scoring functions that are needed and how to weigh them into a single scalar. For example, scoring functions can sometimes not produce the desired chemistry, or show clear exploitation loopholes 48, 49. In this setting, a sufficient element of regularization to a prior policy is useful, as the prior policy has learned a chemical space distribution based on a specified training dataset of choice which acts as a representation of desirable chemistry by examples. It is for this reason that methods like REINVENT have been designed heuristically to contain terms that try to enforce regularization instead of entirely using the reward signal.
@@
To account for this, we make a modification to the average and AUC of the top 10 molecules to ensure they are diverse by an ECFP4 Tanimoto similarity of less than 0.35 to each other. Moreover, we use sphere exclusion diversity (SEDiv@1k) of de novo molecules 47 which measures the proportion of molecules needed to describe chemical space in a random sample of 1,000 molecules, as a proxy for exploration. Note this strategy of measuring diverse hits is of increasing interest 50, 10. This is more representative than the number of unique molecules that could all reside in a close area of chemical space. Lastly, we remove de novo molecules that do not pass a series of filters. This includes a basic chemistry filter (B-CF): logP less than or equal to 4.5, rotatable bond count less than or equal to 7, molecular weight in the range 150 to 650 Da, only contain atoms belonging to the following set $A\in\{C,S,O,N,H,F,Cl,Br\}$, and do not violate the substructure alerts as described in 51. Moreover on the assumption that the training dataset describes desirable chemistry by examples, a target chemistry filter (T-CF): logP and molecular weight within $\mu\pm 4\sigma$ of the training dataset distribution, as well as, removing any molecule that comprises $>10\%$ novel atomic environment bits with respect to the reference molecules, as measured by ECFP4 bits (an example of this is shown in Figure S6) . The combination of the basic and target chemistry filter is denoted (B&T-CF).
@@
Re-evaluating RL performance with these new metrics (Table 2) it can be seen that a smaller proportion of molecules pass the chemistry filters for REINVENT-MolOpt, A2C, PPO, and PPOD, as expected due to decreased regularization terms to optimize efficiency. Despite this, REINVENT-MolOpt achieves the highest score for the average and AUC of the top 10 diverse molecules followed by PPOD. As a proxy for exploration, REINVENT contains the highest sphere exclusion diversity followed by AHC. Interestingly, REINFORCE maintains a high proportion of molecules passing the chemistry filters despite no explicit regularization term, outperforming REINVENT in absolute performance and efficiency. Table 3 and Figure 2b shows the alternative approach of explicitly including these requirements in the reward signal by the addition of the chemistry filters and a diversity filter to each objective. As expected, including these requirements explicitly in the reward function increases the proportion of molecules passing these filters. However, every algorithm drops in performance as measured by the average and AUC top 10 diverse molecules that pass the filters. In this case, this shows that better overall score optimization and efficiency (including those that measure the quality of chemistry) is achieved by implicit regularization to the prior policy. This could be a result of the specific implementation of the chemistry filters in the reward resulting in a score of 0 if they do not pass; however, optimal implementation of reward signal can in itself be considered an art. Therefore, simpler reward signals are usually better, and, achieving desirable chemistry by regularization enables the design of a simpler reward signal. The fine-tuning of the optimization algorithm, reward signals and regularization might be target-dependent and goes beyond the scope of this work, but it is enabled by ACEGEN.
@@
Valid 21.77 $\pm$ 0.03 21.78 $\pm$ 0.02 21.74 $\pm$ 0.05 21.45 $\pm$ 0.02 21.52 $\pm$ 0.15 22.00 $\pm$ 0.02 21.95 $\pm$ 0.02 21.65 $\pm$ 0.06 21.71 $\pm$ 0.03 21.05 $\pm$ 0.14
Top-10 Avg 15.85 $\pm$ 0.10 15.69 $\pm$ 0.11 17.43 $\pm$ 0.17 16.09 $\pm$ 0.10 17.10 $\pm$ 0.19 14.76 $\pm$ 0.09 14.54 $\pm$ 0.10 16.17 $\pm$ 0.16 14.86 $\pm$ 0.08 15.18 $\pm$ 0.24
Top-10 AUC 13.67 $\pm$ 0.07 13.60 $\pm$ 0.08 15.65 $\pm$ 0.14 13.91 $\pm$ 0.08 15.80 $\pm$ 0.14 12.94 $\pm$ 0.07 12.78 $\pm$ 0.08 14.57 $\pm$ 0.14 13.00 $\pm$ 0.06 14.01 $\pm$ 0.18
Unique 22.28 $\pm$ 0.10 22.63 $\pm$ 0.06 13.68 $\pm$ 0.31 22.68 $\pm$ 0.07 10.21 $\pm$ 0.35 22.44 $\pm$ 0.05 22.83 $\pm$ 0.01 16.63 $\pm$ 0.20 22.80 $\pm$ 0.01 14.13 $\pm$ 0.36
B&T-CF 14.34 $\pm$ 0.18 14.74 $\pm$ 0.15 7.00 $\pm$ 0.27 13.77 $\pm$ 0.12 5.71 $\pm$ 0.22 18.44 $\pm$ 0.06 18.59 $\pm$ 0.04 13.62 $\pm$ 0.16 17.72 $\pm$ 0.04 11.55 $\pm$ 0.30
B&T-CF Top-10 Avg (Div) 14.95 $\pm$ 0.12 14.83 $\pm$ 0.13 16.06 $\pm$ 0.16 15.25 $\pm$ 0.10 15.78 $\pm$ 0.15 14.51 $\pm$ 0.09 14.33 $\pm$ 0.08 15.92 $\pm$ 0.13 14.68 $\pm$ 0.08 14.86 $\pm$ 0.22
B&T-CF Top-10 AUC (Div) 12.91 $\pm$ 0.06 12.85 $\pm$ 0.08 14.61 $\pm$ 0.13 13.11 $\pm$ 0.08 14.60 $\pm$ 0.14 12.79 $\pm$ 0.06 12.66 $\pm$ 0.07 14.35 $\pm$ 0.12 12.89 $\pm$ 0.07 13.74 $\pm$ 0.16
B&T-CF Diversity (SEDiv@1k) 17.39 $\pm$ 0.15 18.19 $\pm$ 0.12 10.03 $\pm$ 0.31 17.55 $\pm$ 0.10 7.71 $\pm$ 0.39 16.23 $\pm$ 0.13 17.65 $\pm$ 0.10 7.97 $\pm$ 0.16 17.33 $\pm$ 0.10 7.92 $\pm$ 0.21
Table 3: Comparison of algorithm performance with or without explicit chemistry filters included in the reward signal as a scoring function. MolOpt is the normal benchmark and MolOpt-CF is the benchmark with chemistry filters and a diversity filter applied to each objective. Each algorithm was run 5 times with different seeds, the average score and variance was then summed over the 23 tasks to report the final score and standard deviation of the summed variance.
@@ 3.3 Ablation study of the REINVENT algorithm
One of the most popular RL algorithms used for the fine-tuning of CLMs is REINVENT 4. From a theoretical RL perspective, this algorithm can be viewed as a combination of the REINFORCE algorithm, experience replay, reward shaping and a sequence likelihood penalty term for the loss function. However, the impact of each specific mechanism is unclear. Therefore, using the modular components of ACEGEN, we conducted an ablation study to better understand the impact of reward shaping, experience replay, and sequence likelihood penalty, which are key components of the REINVENT algorithm. We used default REINVENT hyperparameters and measured performance on the MolOpt benchmark for comparison. Regarding performance, Table 4 and Figure 2d indicates that the primary improvement comes from adding experience replay to REINFORCE. While reward shaping does improve over REINFORCE, combining it with experience replay does not further improve performance. However, adding reward shaping that links the agent policy to the prior policy does improve the chemistry quality, measured by the proportion of valid and unique molecules passing the chemistry filters, albeit with a compromise in performance. These results highlight the importance of regularization in practical applications, as discussed in the previous section. In the original REINVENT implementation, the authors also include a loss term to penalize highly likely sequences. Table 4 shows that this primarily increases exploration but negatively affects all other metrics, including reducing the proportion of molecules passing chemistry filters. Furthermore, if used without reward shaping, it results in a complete collapse in learning.
@@
Valid 21.74 $\pm$ 0.04 21.77 $\pm$ 0.03 21.93 $\pm$ 0.02 21.90 $\pm$ 0.02 21.76 $\pm$ 0.02 21.78 $\pm$ 0.02
Top-10 Avg 14.74 $\pm$ 0.09 15.85 $\pm$ 0.10 14.80 $\pm$ 0.10 15.69 $\pm$ 0.10 14.79 $\pm$ 0.11 15.69 $\pm$ 0.11
Top-10 AUC 13.21 $\pm$ 0.07 13.67 $\pm$ 0.07 13.21 $\pm$ 0.07 13.57 $\pm$ 0.08 13.11 $\pm$ 0.07 13.60 $\pm$ 0.08
Unique 22.35 $\pm$ 0.05 22.28 $\pm$ 0.10 22.46 $\pm$ 0.05 22.29 $\pm$ 0.09 22.80 $\pm$ 0.05 22.63 $\pm$ 0.06
B&T-CF 14.15 $\pm$ 0.23 14.34 $\pm$ 0.18 15.31 $\pm$ 0.15 15.08 $\pm$ 0.13 14.83 $\pm$ 0.13 14.74 $\pm$ 0.15
B&T-CF Top-10 Avg (Div) 13.81 $\pm$ 0.09 14.95 $\pm$ 0.12 13.88 $\pm$ 0.09 14.80 $\pm$ 0.11 13.89 $\pm$ 0.12 14.83 $\pm$ 0.13
B&T-CF Top-10 AUC (Div) 12.51 $\pm$ 0.07 12.91 $\pm$ 0.06 12.48 $\pm$ 0.06 12.83 $\pm$ 0.07 12.38 $\pm$ 0.07 12.85 $\pm$ 0.08
B&T-CF Diversity (SEDiv@1k) 17.61 $\pm$ 0.18 17.39 $\pm$ 0.15 17.93 $\pm$ 0.10 17.52 $\pm$ 0.11 18.76 $\pm$ 0.11 18.19 $\pm$ 0.12
Table 4: Ablation study of REINFORCE algorithm on the MolOpt benchmark with and without separate components including experience replay, reward shaping, and a high sequence likelihood penalty. Each algorithm was run 5 times with different seeds, the average score and variance was then summed over the 23 tasks to report the final score and standard deviation of the summed variance. Note that all components together constitute the ACEGEN REINVENT implementation with default hyperparameters.
@@ 3.4 Case Study: De-novo generation in the 5-HT2A
To reflect more realistic situations that may arise in practical drug discovery scenarios, we apply the different RL algorithms implemented on a more challenging objective previously proposed by 42. This objective is designing 5-HT2A receptor ligands selective over the highly related D2 receptor utilizing only structural information. This is an important and relevant challenge as clinically, this off-target profile leads to undesirable extrapyramidal side-effects in the treatment of psychosis 52. To achieve this, this objective aims to minimize the docking score against 5-HT2A (PDB: 6A93), yet maximize the docking score against D2 (PDB: 6CM4), within a budget of 32,000 molecules. Both crystal structures are co-crystallized with Risperidone, reflecting the high degree of binding pocket similarity. In contrast to the proposed objective configuration, we use rDock 53 due to more permissive licensing. We note that although docking scores are generally unreliable as a predictor of binding affinity, they still provide enrichment of selected molecules in virtual screening 54 and have shown benefits over the use of ligand-based prediction of binding affinity 47 which can lead to certain failure modes 48, 49.
@@
Table 5: Performance summary of algorithms on 5-HT2A case study. Including after applying a basic chemistry filter (B-CF) as described previously. B-CF is the fraction of molecules that pass the chemistry filter. Note that the values here are re-normalized based on the scores achieved for a subset of 5-HT2A ligands that display at least 2-fold selectivity over D2 as extracted from ChEMBL31 35. Therefore, a score of 0.91 indicates the best score observed in the known ligand subset, and a score greater than 1.0 indicates that a single molecule achieves both a better (more negative) 5-HT2A docking score and (more positive) D2 docking score than seen anywhere in the known subset.
@@ 3.5 Case study: Scaffold constrained generation
ACEGEN can conduct constrained generation by leveraging PromptSMILES 46 which proposes iterative CLM prompting on SMILES in combination with RL to achieve fine-tune a pre-trained CLMs to the task of constrained generation. Here we demonstrate this application with two experiments: 1) scaffold decoration via known synthetic reactions as proposed with LibINVENT 21, and 2) scaffold decoration followed by scaffold-constrained docking to explore growth vectors inside a binding pocket.
In the first experiment, we compared the performance of RL algorithms on two tasks proposed with LibINVENT 21, both of which conduct scaffold decoration of a piperazine core with an amine linker moiety commonly present in D2 receptor ligands. As in the MolOpt benchmark, we apply a budget of 10,000 molecules. The first task is to optimize the predicted probability of D2 activity as evaluated by a QSAR model. The second task adds a selective reaction filter, rewarding piperazine growth via a Buchwald reaction and amine linker growth via an amide coupling reaction. Table 6 shows successful optimization of both objectives, with AHC performing best on the first task, and REINVENT (either with default or MolOpt parameters) performing best on the second task. However, all algorithms except A2C are able to solve the task with a yield above 90%, and fully satisfy the selective reaction objectives in 80 to 90% of molecules.
@@
D2 Yield 0.977 $\pm$ 0.005 0.989 $\pm$ 0.001 0.987 $\pm$ 0.007 0.991 $\pm$ 0.001 0.354 $\pm$ 0.125 0.942 $\pm$ 0.010 0.952 $\pm$ 0.010
Average score 0.679 $\pm$ 0.008 0.720 $\pm$ 0.011 0.743 $\pm$ 0.020 0.794 $\pm$ 0.002 0.723 $\pm$ 0.043 0.749 $\pm$ 0.003 0.760 $\pm$ 0.006
D2 with Yield 0.972 $\pm$ 0.004 0.992 $\pm$ 0.001 0.988 $\pm$ 0.006 0.990 $\pm$ 0.003 0.555 $\pm$ 0.275 0.950 $\pm$ 0.015 0.947 $\pm$ 0.005
reaction Average score 0.579 $\pm$ 0.003 0.668 $\pm$ 0.003 0.796 $\pm$ 0.007 0.723 $\pm$ 0.003 0.537 $\pm$ 0.149 0.705 $\pm$ 0.028 0.726 $\pm$ 0.013
filters Ratio of satisfied reaction filters 0.896 $\pm$ 0.005 0.753 $\pm$ 0.011 0.921 $\pm$ 0.009 0.809 $\pm$ 0.004 0.352 $\pm$ 0.196 0.830 $\pm$ 0.058 0.858 $\pm$ 0.020
Table 6: Algorithm comparison in combination with PromptSMILES for constrained molecule generation on LibINVENT DRD2 tasks (without and with selective reaction filters). Each algorithm was run 5 times with different seeds, the average value and standard deviation is reported.
In the second experiment, we used PromptSMILES in combination with AHC (based on the superior performance in the 5-HT2A case study) to conduct scaffold decoration in the catalytic site of BACE1, a key therapeutic target for the treatment of Alzheimer’s disease. Here we used the reference structure 4B05 co-crystallized with AZD3839 55 which is a typical BACE1 inhibitor consisting of an amidine core interacting with two catalytic aspartates, and occupation of the P2’, P1 and P3 sub-pockets 56. The bicyclic, amidine-containing core was used as a prompt for PromptSMILES, allowing two growth vectors, one into the P2’ sub-pocket, and one into the P1 and P3 sub-pockets. Molecules were rewarded by minimizing the docking score using substructure constrained docking, as well as maintaining the number of heavy atoms and rotatable bonds in a sensible range (see supporting information F). Figure 4 shows the successful maximization of the reward over the generation of 10,000 molecules (Figure 4a,b), and corresponding minimization of the docking score (Figure 4c). The top 10 molecules identified contain improved docking scores in the range -23.9 to -15.8 compared to the reference molecule re-docked with a docking score of -8.44. Docked poses (shown in Figure S13) show the recovery of aromatic rings in the P1 and P2 sub-pockets. This case study highlights the successful constrained optimization of a complex objective.
Figure 4: Optimization of the multi-objective reward. The average reward and optimization of the underlying docking score are shown. The top 10 de novo molecules are shown by multi-objective reward, with the constrained substructure highlighted in red and the docking score labeled below. For reference, the co-crystal ligand is re-docked with a docking score of -8.44. The docked poses are shown in Figure S13.